OpenStudy (anonymous):

PLEASE HELP. A manufacturer of a brand A jeans has daily production costs of c=0.2x^2-90x+10,700, where C is the total cost (in dollars) and x is the number of jeans produced. How many jeans should be produced each day in order to minimize costs? what is the minimum daily cost?

OpenStudy (ash2326):

@aznbobo We want to minimize the cost. This question of maxima and minima. I'll guide you and you'll solve it. You agree with this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Ok find c'(x)

OpenStudy (anonymous):

minimum daily cost would be when you produce nothing...

OpenStudy (anonymous):

@aznbobo did you got my point???

OpenStudy (ash2326):

@sheg We'll just find out that x=0 won't give us the lowest cost

OpenStudy (ash2326):

@aznbobo Did you find c'(x)

OpenStudy (anonymous):

i do not know how to do that. >.>

OpenStudy (anonymous):

@ash2326 plug in x=0 and see what you get it is fixed cost that incur to the company when the company produces virtually nothing

OpenStudy (anonymous):

when i plug in 0 for x i get... 10700

OpenStudy (ash2326):

@aznbobo that's not right. Do you know differentiation?

OpenStudy (anonymous):

now if you are talking of finding out the minimum cost by finding the total number of jeans produced so in this case we can use the Khun-Tucker condition which states that if the first derivative of any equation is equal to zero and the second derivative is positive then we can get the minimum cost @ash2326

OpenStudy (anonymous):

so differentiate the given equation

OpenStudy (ash2326):

@aznbobo Do you know parabolas?

OpenStudy (anonymous):

Kuhn-Tucker Condition*

OpenStudy (anonymous):

yes parabolas

OpenStudy (anonymous):

@aznbobo you just need to differentitate the given equation with respect to x

OpenStudy (ash2326):

@aznbobo how's the graph of \[c(x)=0.2x^2-90x+10700\] Does it have concavity upwards or downwards?

OpenStudy (anonymous):

upwards.

OpenStudy (anonymous):

on diff. you will get 0.4x - 90 = 0 again diff. wrt x 0.4 that is a positive quantity so the above equation fulfills the Kuhn-Tucker Condition so above equation i.e., 0.4x - 90 = 0 0.4x = 90 x = 225 so number of jeans produced would be 225 and by plugging in x = 225 you can calculte the minimum cost

OpenStudy (ash2326):

Good you know this function is our cost function. The cost will be minimum at the vertex, can you find the vertex of this graph?

OpenStudy (anonymous):

umm i think (225,275)?

OpenStudy (ash2326):

How did you find it? Can you explain a little?

OpenStudy (anonymous):

i graphed in on the calculator..

OpenStudy (ash2326):

You know standard form of a parabola is \[y=a(x-h)^2+k^2\] Where h,k is the vertex here \[c(x)=a(x-h)^2+k^2\] We need to convert the equation in this form and we'll get the minimum cost as k and no. of jeans as h. Do you understand this?

OpenStudy (anonymous):

so k is already given at 10700, right? h would be... 90?

OpenStudy (ash2326):

No it's not easy, we need to convert in that form We have \[c(x)=0.2x^2-90x+10700\] We need it an a form \[c(x)=\underline{a(x-h)^2}+k^2\] Notice the underlined term here we need to have a perfect square term (x-h)^2 do you understand the difference?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Great, ok let's get to work now

OpenStudy (ash2326):

We have \[c(x)=0.2x^2-90x+10700\] now this can be written as \[c(x)=0.2(x^2-450x)+10700\] now we have \[x^2-450\] We need to convert this to a square term, Can you do that?

OpenStudy (anonymous):

x=\[15\sqrt{2} \] & \[-15\sqrt{2} \]?

OpenStudy (ash2326):

@aznbobo We need to convert this to a perfect square term \[x^2-450x \longrightarrow(x-a)^2\] Can you do it? If you can't I'll help you:)

OpenStudy (anonymous):

mmm i can you help me please? :]

OpenStudy (ash2326):

Ok \[x^2-450x\] Whenever you need to convert into a perfect square, add and subtract the square of half the coefficient of x here coefficient of x is 450, so half of it is 225, We'll add and subtract the square of this, so we have \[x^2-450x+225^2-225^2\] so We get \[(x-225)^2-225^2\] so We'll substitute this for x^2-450x \[x^2-450x\longrightarrow (x-225)^2-225^2\] Do you understand this?

OpenStudy (anonymous):

okay

OpenStudy (ash2326):

Now we have \[c(x)=0.2((x-225)^2-225^2)+10700\] It can be written as \[c(x)=0.2(x-225)^2-0.2 \times 225^2+10700\] \[c(x)=0.2(x-225)^2-0.2 \times 50625+10700\] \[c(x)=0.2(x-225)^2- 10125+10700\] Finally we get \[c(x)=0.2(x-225)^2+ 575\] So we have the form of \[c(x)=a(x-h)^2+k\] here k is the minimum cost and h is the no. of jeans for minimum cost, the vertex is h,k here it's h=225 and k=575 so no. of jeans=225 minimum cost= 575

OpenStudy (ash2326):

@aznbobo Do you understand this?

OpenStudy (anonymous):

@ash2326 thats what i calculated using the Kuhn_Tucker Condition